Liquidating Partnerships

Prior to liquidating their partnership, Todd and Dunn had capital accounts of $66,000 and $101,000, respectively. Prior to liquidation, the partnership had no cash assets other than what was realized from the sale of assets. These partnership assets were sold for $149,000. The partnership had $8,000 of liabilities. Todd and Dunn share income and losses equally.

Determine the amount received by Todd as a final distribution from liquidation of the partnership.
$

Prior to liquidating their partnership, Pepper and Reynell had capital accounts of $8,000 and $32,000, respectively. The partnership assets were sold for $16,000. The partnership had no liabilities. Pepper and Reynell share income and losses equally.

Required:

a. Determine the amount of Pepper's deficiency.
$

b. Determine the amount distributed to Reynell, assuming Pepper is unable to satisfy the deficiency.
$
Answer #1
1
Total Capital 167000 =66000+101000
Book value of assets 175000 =167000+8000
Loss on sale of assets 26000 =175000-149000
Todd Capital 66000
Less: Share of loss -13000 =26000*50%
Amount received by Todd 53000
2
Loss on sale of assets 24000 =8000+32000-16000
a
Pepper, capital 8000
Less: Share of loss -12000 =24000*50%
Pepper's deficiency 4000
b
Reynell Capital 32000
Less: Share of loss -12000
Less: Pepper's deficiency -4000
Amount distributed to Reynell 16000